OpenStudy (deaththekid975):

A jar contains 12 nickels, 12 dimes, and 12 quarters. A coin is chosen at random from the jar. What is the probability that the coin chosen is a quarter? A. 1/36 B. 1/12 C. 1/4 D. 1/3

OpenStudy (thatemokid66):

okay so add them all: 12+12+12=36 there are 12 quarters so 12/36 = 1/3

OpenStudy (thatemokid66):

the answer is D

OpenStudy (madiimay4music):

12*3 = 36 tho

jabez177 (jabez177):

@ThatEmoKid66, Hello and Welcome to OpenStudy! :) You seem new here and more than willing to help which is mighty fine since we need all the help we can get :) but you might want to take some time and read this site's code of conduct! http://openstudy.com/code-of-conduct This site has a policy for "Give Help, Not Answers" violations of such policies could potentially result in an account suspension or ban. It's great to see another helper among us, so it would be sad if you got in trouble for giving out answers. Instead, you can try guiding the asker through their question to help them understand it. Hope you'll have a nice time here on OpenStudy, and I'll see ya around :) Have a Nice Day!

OpenStudy (madiimay4music):

At least you weren't mean about it.

OpenStudy (thatemokid66):

@jabez177 Sorry :(